# PaperRegister: Boosting Flexible-grained Paper Search via Hierarchical Register Indexing

### v1

@@ Introduction
We first use LitSearch (Ajith et al. 2024), mainly containing coarse-grained queries. And then we build a new dataset, Flexible-grained Search, covering queries across various granularity and including data for test, development, and training. The detailed statistical information of dataset is shown in Appendix D. And the metrics used and reported in experiments are R@5 (recall@5) and R@10 (recall@10).
(1) LitSearch (Ajith et al. 2024). By using GPT-4 to rewrite citations in papers and asking authors to write, this dataset constructs 597 paper search queries, mainly involving coarse-grained topics like “Where can I find research on evaluating consistency in generated summaries”.
(2) Flexible-grained Search. We collect 4,200 papers in machine learning field as corpus and perform the offline hierarchical indexing. To build test data across various granularity, we pick hierarchical register of 2,100 papers, employ Qwen3-32B to find original paper texts related to each content in register, then use Qwen3-32B to generate queries based on these texts, obtaining 5,644 test data including 3 parts: general-granularity (F.g.Search-1), fine-granularity (F.g.Search-2), and very fine-granularity (F.g.Search-3). The format of each test data is like $\{q_{j},p_{j}\}$. To build training and development data, we use register of another 2,100 paper and directly generate queries for contents in each paper register, where each query is with a golden view, the format of training and development data is like $\{q_{j},v_{j},p_{j}\}$. Then we do random splitting and get 13,824 training data and 695 development data, the training data are used in view recognizer training part as illustrated in the Formula 9 and development data are used for analysis in the Table 3.
@@
We select baselines from commonly used or recent methods suitable for paper search tasks.
(1) Direct Matching. These methods directly use paper titles, abstracts or total text of papers to construct index of corpus, and retrieval by matching query with these contents.
(2) Query-paraphrasing Methods. These methods aim to improve paper‐searching performance by paraphrasing the input query, including: Rewriting (Ma et al. 2023), which uses a LLM to rewrite original query. HyDE (Gao et al. 2023), which uses a LLM to generate a fake document based on input query and retrieves real documents by this fake document. CSQE (Lei et al. 2024), which initially retrieves several documents and then uses a LLM to expand original query based on these initially-retrieved documents. We use Qwen3-32B as the LLM when conducting these baselines.
(3) Paper-splitting Methods. These methods split the original paper into multiple short parts, then calculate the similarity between the query and each short part, and finally integrate all the similarity to determine the overall similarity of query and original paper. In experiments, we employ four kinds of settings, including splitting the original paper to the fixed length of 512-token chunk or by the raw paragraph, taking the average similarity or the maximum similarity of all parts as the final score, represented as $\text{Chunk}_{avg}$, $\text{Chunk}_{max}$, $\text{Paragraph}_{avg}$, and $\text{Paragraph}_{max}$, respectively.
Method BM25-based Matching DPR-based Matching
@@
Table 2: Ablation results for the hierarchical register indexing. The table shows that hierarchical index tree are important for flexible-grained paper search and different layers in the hierarchical register schema serve queries at different granularity.
Implementation Details.
In the offline stage, we employ Qwen3-32B (Qwen 2025) as the large language model in process of fine-grained content extracting and bottom-up content aggregating, and deploy it as API by vllm [^1] https://pypi.org/project/vllm/ on two A-100 80G GPU for convenience. In the online stage, we set $K$ as 5 and $M$ as 5 or 10, use Qwen3-0.6B (Qwen 2025) as the base model of view recognizer, employ a prefix tree-based restricted decoding strategy (Tang et al. 2024) in the view identifying process to prevent the model outputting irrelevant token, and conduct the view-based matching process via rank-bm25222https://pypi.org/project/rank-bm25/ for BM25 and gte-Qwen2-7B-instruct (Li et al. 2023) for DPR. For the view recognizer training, we use TRL [^3] https://github.com/huggingface/trl framework to conduct SFT and GRPO. We set train epoch as 5 in SFT, $G$ as 5 and train epoch as 2 in GRPO, and keep all other parameters as default value in TRL. When conducting baselines, we keep the same settings as above.
Overall Results
@@
Results compared with baselines are shown in the Table 1, there are two main conclusions based on the results:
@@
(1) PaperRegister is a powerful solution to addressing flexible-grained paper search tasks. As shown in the Table 1, PaperRegister demonstrates excellent performance in paper search tasks at various granularity, outperforming all baselines in both BM25-based matching and DPR-based matching settings, on both recall@5 and recall@10 metrics. For example, in F.g.Search-3, under the DPR-based matching setting, PaperRegister achieves a recall@5 score of 80.8, while using abstract-based index yields only 58.2, where PaperRegister can achieve a performance improvement of 22.6. All in all, PaperRegister can improve a lot compared with various previous paper search methods.
Figure 4: Paper search performance of the PaperRegister with different view recognizer. The figure shows strong recognizer is with obvious positive impact on overall system.
@@
(2) Compared to traditional methods, the advantages of PaperRegister become more pronounced at these finer granular queries. According to experimental setting, from LitSearch to F.g.Search-1 to F.g.Search-2 to F.g.Search-3, the granularity of the paper search query becomes increasingly finer. And the experimental results show that the performance improvement of PaperRegister over the method using abstract-based index becomes larger. For example, under the DPR-based matching setting, PaperRegister achieves recall@5 scores of 81.0, 84.1, 79.9, and 80.8 on the four datasets, respectively, while using the abstract-based index yields scores of 77.7, 69.4, 62.1, and 58.2. Here, PaperRegister delivers performance improvements of 3.3, 14.7, 17.8, and 22.6, respectively. Therefore, PaperRegister exhibits more improvements over baseline on finer-grained queries, which can to some extent validate rationality of this work’s motivation, that traditional paper search via abstract-based index struggles to handle fine-grained queries.
Ablation Results
@@
To validate the importance of hierarchical register indexing, we employ ablation experiments by simplifying the hierarchical register schema. As shown in Table 2, where “w/ only layer-1” means retaining only the coarsest-grained nodes, “w/ only layer-2” means retaining only the medium-grained nodes, and “w/ only layer-3” means retaining only the finest-grained nodes. Based on these simplified register schema for experiments, we get the following conclusions:
@@
(1) Hierarchical register indexing is important for ensuring the effectiveness of paper search. As shown in Table 2, compared to using the complete hierarchical register schema, using a register schema composed of nodes from any single layer leads to obvious performance drop. For example, recall@5 of LitSearch under BM25-based matching drops from 69.7 to 64.5, 66.8, 64.6 for “w/ only layer-1”, “w/ only layer-2”, “w/ only layer-3”, respectively. And other datasets also show similar situation. Therefore, results strongly prove importance of hierarchical register indexing.
@@
(2) Different layers in the hierarchical register schema serve queries at different granularity. For example, under “w/ only layer-1” setting, the DPR-based matching and recall@5 metric of F.g.Search-1 drops from 84.1 to 82.5 (1.6 decrease), while F.g.Search-1 is from 80.8 to 68.1 (12.7 decrease), which is significantly larger than 1.6. Conversely, under “w/ only layer-3” setting, F.g.Search-1 drops from 84.1 to 77.8 (6.3 decrease), while F.g.Search-3 is from 80.8 to 80.4 (0.4 decrease), which is much smaller than 6.3. Therefore, nodes from different layers play specific roles in handling queries of different granularity, further strongly demonstrating necessity of hierarchical register indexing.
@@
Detailed Analysis
In this section, we conduct detailed analysis including: the effect of view recognizer, the compatibility of PaperRegister with complex PaSa framework (He et al. 2025), and an analysis of the online search efficiency. Due to the space constraint, we only report the performance that is based on DPR-based matching for analysis in this section.
Effect of View Recognizer and Training.
To analyze the role of view recognizer in PaperRegister and necessity of our training, we first examine relationship between recognizer capability and PaperRegister performance in Figure 4, then compare with Qwen3-32B and conduct ablation in Table 3.
As shown in Figure 4, where “Bad Recognizer”,“Random Recognizer”,“Weak Recognizer” represent the completely incorrect, randomly predicting, and weak-performing view recognizers, respectively, the curve demonstrates a clear positive correlation between the capability of view recognizer and final performance of PaperRegister. Therefore, building an accurate view recognizer is a key factor for the PaperRegister to achieve high-performance paper search.
@@
As shown in Table 3, We compare accuracy and latency of view identifying between the view recognizer in PaperRegister and Qwen3-32B (enable thinking) under zero-shot and few-shot settings. The results show the view recognizer in PaperRegister is with better accuracy and latency than powerful Qwen3-32B. Furthermore, we only keep SFT for the view recognizer training, or replace the hierarchical reward with direct 0-1 reward in GRPO training. Results show that the accuracy decreases to some extent. Therefore, training a small-scale language model with SFT and then hierarchical-reward GRPO is an effective approach to obtain the view recognizer with both high performance and low latency.
Compatibility with Complex PaSa Framework.
@@
Considering that PaperRegister is mainly a retrieval module, and given that some recent complex paper search frameworks such as PaSa (He et al. 2025) incorporate various modules like rewriting, retrieval, iteration, and filtering, in order to explore whether PaperRegister can be compatible with such complex frameworks and further enhance performance, we replace the original retrieval module in PaSa with PaperRegister. As shown in Figure 5, PaperRegister can further improve the performance of PaSa across paper search tasks at various granularity. Therefore, PaperRegister can be effectively adapted as a retrieval module into complex paper search frameworks and further enhance the capability.
@@
Figure 5: Performance of adding PaperRegister into the PaSa framework. The figure shows that PaperRegister can greatly cooperate with extra complex modules in PaSa framework.
Analysis for Online Search Efficiency.
@@
Considering that the efficiency of online paper search system is a crucial aspect of its usability in practical scenarios, we compare the online search efficiency of PaperRegister with multiple baselines. As shown in the Table 4, the online latency of these baselines is significantly higher than that of PaperRegister, which limits the practical applicability of these methods in real-world scenarios. In contrast, PaperRegister demonstrates the acceptable online latency, proving its good potential for application in the real-world paper search.
Related Work
@@
To achieve accurate paper search, previous related works mainly focus on enhancing the input query, with less attention given to the index of paper corpus. Some works improve paper search performance by rewriting the original query, such as using powerful large language models for rewriting based on some manually designed instructions (Ma et al. 2023; Anand et al. 2023), generating pseudo-documents to replace original query for retrieval real documents (Gao et al. 2023; Li et al. 2024a), or employing several powerful models as agent system to conduct multi-round rewriting and expansion (He et al. 2025; Ren et al. 2025). Another kind of works expands the input query through pseudo-relevance feedback, such as augmenting the query based on initially retrieved documents (Lei et al. 2024; Li et al. 2025a), or extracting several keywords from the corpus to enrich the information of original query (Kang et al. 2024; Zhang et al. 2025). Although these works can improve performance in regular paper search, they fail to address flexible-grained queries because they do not solve the flaws in the index of paper corpus. Recently, several works split original paper by fine-grained fields and then construct multiple-field indexes for retrieval (Sotaro et al. 2024; Li et al. 2025b; Shi et al. 2025; Chen et al. 2025). However, these works still cannot effectively handle flexible-grained queries because they only focus on a single granularity level.
Compared with above related works, the largest advantage of PaperRegister is hierarchical register indexing and conducting paper search with hierarchical index tree, which can effectively handle flexible-grained paper search.
Method Online Paper Search Latency (s) $\downarrow$
Rewriting 9.3
HyDE 20.7
CSQE 33.5
$\text{Chunk}_{max}$ 5.4
$\text{Paragraph}_{max}$ 4.2
PaperRegister (Ours) 2.5
@@
Table 4: Comparison of online latency among PaperRegister and several baselines. The table show that PaperRegister is with better real-world applicability with less latency.
Conclusion
@@
In this work, noticed limitation of existing methods in paper search tasks, we propose PaperRegister, boosting flexible-grained paper search via hierarchical register indexing. In the offline stage, PaperRegister constructs a hierarchical index tree by fine-grained content extracting and bottom-up content aggregating. In the online stage, PaperRegister first identify views of input query, then matches the query against corresponding indexes from the hierarchical index tree. Furthermore, we construct the view recognizer by applying supervised finetuning and hierarchical-reward group relative policy optimization to a small-scale language model. Experiments on extensive flexible-grained paper search tasks demonstrate that PaperRegister is an effective solution, which reaches the SOTA performance, with improvement being more pronounced as query granularity finer. Therefore, this work offers a promising direction for developing more powerful paper search systems in future.
Appendix A
This part is about detailed illustration for the hierarchical register schema.
Figure 6: The first kind of hierarchical register schema.
Figure 7: The second kind of hierarchical register schema.
Figure 8: The third kind of hierarchical register schema.
Figure 9: The fourth kind of hierarchical register schema.
Figure 10: The fifth kind of hierarchical register schema.

### v2

@@ 4.1 Experimental Settings
To conduct experiments for flexible-grained paper search, we first use LitSearch ajith2024litsearch, which mainly contains coarse-grained queries, and then we build Flexible-grained Search, a new dataset covering queries across various granularity, including data for test, development and training. The metrics reported in experiments are R@5 (recall@5) and R@10 (recall@10). And detailed statistical information is in the Appendix D.
(1) LitSearch ajith2024litsearch. By using GPT-4 to rewrite citations in papers and asking authors to write, this dataset constructs 597 paper search queries, mainly involving coarse-grained topics like “Where can I find some researches on evaluating consistency in generated summaries”.
(2) Flexible-grained Search. We collect 4,200 papers in the arXiv platform as corpus and perform the offline hierarchical indexing. To build training and development data, we pick 2,100 paper registers, and generate query based on each node content in each register by LLM, where the node path is golden view of query. Then we randomly split, getting 13,824 training data and 695 development data, training data are used in view recognizer training, and development data are for analysis in Table 3. The format of training and development data is $\{q_{j},v_{j},p_{j}\}$. To build test data across various granularity, we pick the other 2,100 paper registers. And in order to prevent data leakage, we do not directly use node content to generate query. Instead, we employ LLM to find original paper text related to each node, then generate queries based on these text, obtaining 5,644 test data. Thanks to the hierarchical feature of register, test data cover: F.g.Search-1 (general granularity), F.g.Search-2 (fine granularity), and F.g.Search-3 (very fine granularity). Format of each test data is $\{q_{j},p_{j}\}$, and LLM in this section is Qwen3-32B qwen3.
@@
We select baselines from commonly used or advanced methods suitable for paper search tasks. (1) Direct Matching. These methods directly use the paper title or abstract to construct corpus index, and then perform retrieval by matching query with these contents. In addition, to demonstrate that the effectiveness of PaperRegister does not simply stem from using total paper text, we include a method that constructs the corpus index using the total paper text into this category of baselines. (2) Query Enhancing. These methods aim to improve paper search performance by paraphrasing the input query, including: Rewriting ma2023query, which uses a LLM to rewrite original query. HyDE gao2023precise, which uses a LLM to generate a fake document based on input query and retrieves real documents by this fake document. CSQE lei2024corpus, which initially retrieves several documents and then uses a LLM to expand original query based on these initially-retrieved documents. We use Qwen3-32B as LLM for these baselines. (3) Multi-field Indexing. These methods split original document into multiple parts, get flat multi-field index, then calculate similarity by selected indexes and integrate to determine overall relevance li2025multifield. In experiments, due to the lack of a universal multiple-filed partitioning for all kinds of paper, we employ four settings, including splitting the original paper to fixed length of 512-token chunk or by raw paragraph, taking average similarity or maximum similarity of all parts as final score, represented as $\text{Chunk}_{avg}$, $\text{Chunk}_{max}$, $\text{Paragraph}_{avg}$, and $\text{Paragraph}_{max}$, respectively.
Implementation Details.
In the offline stage, we employ Qwen3-32B qwen3 as the large language model in process of fine-grained content extracting and bottom-up content aggregating, and deploy it as API by vllm [^1] https://pypi.org/project/vllm/ on two A-100 80G GPU for convenience. In the online stage, we set $K$ as 5 and $M$ as 5 or 10, use Qwen3-0.6B qwen3 as the base model of view recognizer, employ a prefix tree-based restricted decoding strategy tang2024selfretrieval in view identifying process to prevent the model outputting irrelevant token, and conduct view-based matching process via rank-bm25222https://pypi.org/project/rank-bm25/ for BM25 and gte-Qwen2-7B-instruct li2023towards for DPR. For the view recognizer training, we use TRL [^3] https://github.com/huggingface/trl framework to conduct SFT and GRPO, with train epoch as 5 in SFT, $G$ as 5 and train epoch as 2 in GRPO, and all other parameters as default value.
Method BM25-based Paper Search DPR-based Paper Search
@@
Table 2: Ablation for hierarchical register indexing. Results show that hierarchical index tree is important for flexible-grained paper search and different layers in hierarchical register schema serve queries at different granularity.
4.2 Overall Results
@@
Results compared with baselines are shown in the Table 1, there are two main conclusions:
@@
(1) PaperRegister is an effective system to addressing flexible-grained paper search. As shown in the Table 1, PaperRegister demonstrates excellent performance in paper search tasks at various granularity, outperforming all baselines in both BM25-based matching and DPR-based matching settings, on both recall@5 and recall@10 metrics. For example, in F.g.Search-3, under the DPR-based matching setting, PaperRegister achieves a recall@5 score of 80.8, while using abstract-based index yields only 58.2, where PaperRegister can achieve a performance improvement of 22.6. All in all, PaperRegister can improve a lot compared with various previous paper search methods.
@@
(2) Compared to traditional methods, advantage of PaperRegister become more pronounced at finer granular queries. According to experimental setting, from LitSearch to F.g.Search-1 to F.g.Search-2 to F.g.Search-3, the granularity of query becomes increasingly finer. Experimental results show performance improvement of PaperRegister over abstract-based method becomes larger. For example, under the DPR-based matching setting, PaperRegister achieves recall@5 scores of 81.0, 84.1, 79.9, and 80.8 on the four datasets, respectively, while using the abstract-based index yields scores of 77.7, 69.4, 62.1, and 58.2. Here, PaperRegister delivers performance improvements of 3.3, 14.7, 17.8, and 22.6, respectively. This can to some extent validate rationality of our motivation, traditional paper search methods via abstract-based index struggles to handle fine-grained queries.
4.3 Ablation Results
@@
To validate importance of hierarchical register indexing, we employ ablation by simplifying register schema. As shown in Table 2, where “w/ only layer-1” means retaining only the coarsest-grained nodes, “w/ only layer-2” means retaining only the medium-grained nodes, and “w/ only layer-3” means retaining only the finest-grained nodes. Base on the ablation results, we get the following conclusions:
@@
(1) Hierarchical register indexing is important for ensuring accurate paper search. Compared to using complete hierarchical register schema, using a schema composed of nodes from any single layer leads to obvious performance drop. For example, recall@5 of LitSearch under BM25-based matching drops from 69.7 to 64.5, 66.8, 64.6 for “w/ only layer-1”, “w/ only layer-2”, “w/ only layer-3”, respectively. And other datasets also show similar situation. Therefore, results strongly prove the importance of hierarchical register indexing.
Figure 4: Performance of PaperRegister with different view recognizer. The figure shows a strong recognizer is with obvious positive impact on the overall system.
@@
(2) Different layers in hierarchical register schema serve queries at different granularity. For example, under “w/ only layer-1” setting, the DPR-based matching and recall@5 metric of F.g.Search-1 drops from 84.1 to 82.5 (1.6 decrease), while F.g.Search-1 is from 80.8 to 68.1 (12.7 decrease), which is significantly larger than 1.6. Conversely, under “w/ only layer-3” setting, F.g.Search-1 drops from 84.1 to 77.8 (6.3 decrease), while F.g.Search-3 is from 80.8 to 80.4 (0.4 decrease), which is much smaller than 6.3. Therefore, nodes of different layers play specific roles in handling queries of different granularity, further strongly proving necessity of hierarchical register indexing.
4.4 Detailed Analysis
In this section, we conduct detailed analysis to explore the compatibility and real-world practicality of PaperRegister. Due to the space constraint, we only report DPR-based performance for analysis.
Effect of View Recognizer and Training.
In order to strictly analyze the role of view recognizer in PaperRegister and the necessity of our training, we first examine relationship between the capability of view recognizer and the final performance of PaperRegister in Figure 4, then compare with Qwen3-32B and conduct ablation in Table 3. The detailed process and findings are in following:
As shown in Figure 4, where “Bad Recognizer”,“Random Recognizer”,“Weak Recognizer” represent the completely incorrect, randomly predicting, and weak-performing view recognizers, respectively, the curve demonstrates a clear positive correlation between the capability of view recognizer and final performance of PaperRegister. Therefore, building an accurate view recognizer is a key factor for the PaperRegister system to achieve high performance in complex paper search.
@@ 4.4 Detailed Analysis
As shown in Table 3, We compare accuracy and latency of view identifying between the view recognizer in PaperRegister and Qwen3-32B (enable thinking) under zero-shot and few-shot settings. The results show the view recognizer in PaperRegister is with better accuracy and latency than powerful Qwen3-32B. Furthermore, we only keep SFT for the view recognizer training, or replace the hierarchical reward with direct 0-1 reward in GRPO training. Results show that accuracy decreases to some extent. Therefore, hierarchical-reward GRPO is effective to enhance the view recognizer achieving both high performance and low latency.
Compatibility with PaSa.
@@
Given that some complex information retrieval frameworks such as PaSa he2025pasa incorporate various modules like rewriting, retrieval, iteration, and filtering, to explore whether PaperRegister can be compatible with such complex frameworks and further enhance paper search performance, we conduct experiment by replacing the original search module in PaSa with PaperRegister. As shown in Figure 5, PaperRegister can further improve performance of PaSa on paper search tasks across various granularity. Therefore, PaperRegister can be effectively adapted as a search system into complex frameworks and further enhance the paper search capability.
@@
Figure 5: Performance of combining PaperRegister into PaSa framework. The figure shows that PaperRegister can greatly cooperate with complex modules in PaSa.
Online Search Efficiency.
@@
Considering that the efficiency of online paper search system is a very crucial aspect for ensuring its usability in real-world scenarios, we compare the online search efficiency of PaperRegister with multiple baselines. As shown in the Table 4, the online latency of these baselines is significantly higher than that of PaperRegister, which limits the practical applicability of these methods in real-world scenarios. In contrast, PaperRegister demonstrates relatively acceptable online latency, proving it has a good potential for application in the real-world paper search tasks.
Reducing Indexing Time.
Although PaperRegister achieves excellent performance, it requires a relatively long time for offline indexing. Therefore, we explore whether it can sacrifice a certain amount of performance to reduce time consumption. As shown in Table 5, we replace original two-step registering (extracting and aggregating) with directly extracting all contents. Results show this can significantly reduce the offline indexing time with acceptable performance degradation, which provides an alternative approach allowing a trade-off based on specific requirements in real-world application.
5 Related Work
Query Enhancing.
@@ 5 Related Work
Previous works mainly improve paper search by query enhancing, such as using LLM rewriting ma2023query; anand2023context, generating pseudo-documents to replace original query gao2023precise; li2024automir, employing powerful models as agent system for multi-round expansion he2025pasa; ren2025towards, augmenting via initially retrieved documents lei2024corpus; li2025pseudo, or extracting keywords from corpus to enrich query kang2024taxonomy; zhang2025scientific. Although these methods can improve regular paper search, they fail to address flexible-grained paper search because do not solve fundamental flaw of abstract-based index.
Multi-filed Indexing.
Recently, several studies improve by splitting documents into multiple parts to build a flat, multi-field index takeshita-etal-2024-gengo; li2025multifield; shi2025hypercube; chen2025enrichindex. However, these works are mainly limited to a single granularity level, still struggling to adapt to paper search that demands varying degrees of granularity. In contrast, PaperRegister fundamentally differs by introducing a hierarchical register index. Tree-structured organization contains information across various granularity, enabling system to conduct matching at the appropriate levels, thereby effectively supporting flexible-grained paper search.
Method Avg. R@5 Avg. R@10 Online Latency (s) $\downarrow$
Rewriting 64.7 70.5 9.3
HyDE 64.6 70.9 20.7
CSQE 64.8 69.8 33.5
$\text{Chunk}_{max}$ 77.0 81.4 5.4
PaperRegister (Ours) 81.5 84.9 2.5
@@
Table 4: Comparison of online latency. PaperRegister is with less latency for better real-world applicability.
Method Avg. R@5 Avg. R@10 Indexing Time (h) $\downarrow$
Abstract 66.9 72.1 /
PaperRegister (Ours) 81.5 84.9 47.6
w/o two-step registering 76.1 78.9 14.5
Table 5: PaperRegister can support sacrificing certain performance to alleviate offline time consumption.
6 Conclusion
@@
In this work, we propose PaperRegister, which can support paper search queries at flexible granularity. Specially, PaperRegister offline constructs a hierarchical index tree and online adaptively retrieve based on this tree. Furthermore, we design a powerful and low-latency view recognizer by applying supervised finetuning and hierarchical-reward group relative policy optimization. Experiments on flexible-grained paper search tasks demonstrate PaperRegister is an effective solution, with improvement being more pronounced as granularity finer. This work offers a promising direction for developing more powerful paper search systems in future.
7 Limitations
PaperRegister transforms raw paper into hierarchical paper register, thereby achieving strong performance in flexible-grained paper search. The main limitations are on the consumption associated with LLM utilization and risks stemming from the inherent limitations of LLM. Firstly, using LLM for offline hierarchical indexing requires substantial computational and storage resources. When the paper corpus is large, this may place high demands on computer hardware. Secondly, our experiments are conducted only on normal paper corpus, where LLM could effectively accomplish the specified task. We have not yet do validate under extreme conditions, such as with incomplete papers or papers from niche domains, primarily due to a lack of relevant experimental data. Therefore, mitigating computational resource consumption of PaperRegister, as well as verifying and improving effectiveness and robustness across a broader range of paper, remain important directions for the future work.
Appendix A Detailed illustration for hierarchical register schema